# Hanke-Raus heuristic rule for iteratively regularized stochastic gradient descent

### v1

@@ 3. Iteratively regularized Stochastic gradient descent in Hilbert spaces
This section covers the convergence analysis of IRSGD method under certain assumptions. Precisely, this section is organised as follows. The subsection 3.1 aims to discuss the assumptions required for studying the convergence analysis. In subsection 3.2, we discuss a heuristic rule based on the modified discrepancy principle. Convergence of non-noisy iterates is discussed in subsection 3.3. Lastly, subsection 3.4 aims to demonstrate the convergence analysis of noisy iterates.
We recall that the IRSGD method can be expressed as
((3.1)) $u_{k+1}^{\delta}=u_{k}^{\delta}-\omega_{k}F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0}),\hskip 14.22636ptk\geq 0,$
with $\text{with}\ 0\leq\lambda_{k}\leq\lambda_{\text{max}}<\frac{1}{2}$, where $F^{\prime}_{i_{k}}$ is the Frećhet derivative of $F_{i_{k}}$, $u_{0}^{\delta}$ is the initial guess of the exact solution, $\omega_{k}$ is the step size and the index $i_{k}$ is drawn randomly from the index set $\{0,1,\ldots,P-1\}$ at each iteration step.
@@ 3.1. Assumptions
In our work, we need the following assumptions.
@@
For $i\in\{0,1,\ldots,P-1\}$ and $\rho>0$, there holds:
@@
    The operator $F:U\rightarrow Y^{P}$ is continuous.
@@
    There exists a constant $L>0$ such that
    ((3.2)) $\|F_{i}^{\prime}(u)\|\leq L\ \text{for all}\ u\in B_{\rho+c(\rho)}(u_{0}),$
    where $L\leq\frac{1}{2\omega}$ (see Lemma 3.1 for the meaning of $c(\rho)$ and $\omega$). Moreover there is a constant $10<\eta<$ 1 such that
    ((3.3)) $\|F_{i}(\overline{u})-F_{i}(u)-F_{i}^{\prime}(u)(\overline{u}-u)\|\leq\eta\|F_{i}(\overline{u})-F_{i}(u)\|$
@@
Motivated by [9], we consider the following heuristic rule for IRSGD in Hilbert spaces.
Rule 1.
Let $a\geq 1$ be a fixed number and let
$\Psi(k,y^{\delta}):=(k+a)\|F(u^{\delta}_{k})-y^{\delta}\|^{2}.$
We define $k_{*}:=k_{*}(y^{\delta})$ to be an integer such that
@@
where $k_{\infty}:=k_{\infty}(y^{\delta})$ is the largest integer such that $u^{\delta}_{k}\in\mathfrak{D}(F)$ for all $0\leq k<k_{\infty}$.
@@
Rule 1 is straightforward to implement in practice, as detailed in Section 5. During the iteration process, the value of $\Psi(k,y^{\delta})$ is monitored as a function of $k$. After a sufficient number of iterations, the process is terminated and $k_{*}$ is selected that minimizes $\Psi(k,y^{\delta})$. However, an important challenge associated with the nonlinear Landweber iteration [11] lies in determining an appropriate upper bound on the number of iterations. This is particularly critical due to its local convergence properties, which can lead to premature stopping or inaccurate identification of $k_{*}$. Therefore, the limit on the upper bound must be sufficiently large to avoid inaccurately estimating $k_{*}$ at the first local minimum, which is often not the global minimum [11]. To prevent $k_{*}$ from becoming too small, we have performed several numerical experiments on different values of $a$ in Rule 1.
Note that in order to prove Lemma 3.1 for IRSGD, it is essential to assume that $a\geq 1$. This Lemma will be further required to show the convergence of the scheme. Using the integer $k_{*}:=k_{*}(y^{\delta})$ determined by Rule 1, we call $u^{\delta}_{k_{*}(y^{\delta})}$ as an approximate solution.
A natural question arises concerning a family of noisy data $\{y^{\delta}\}$, where $y^{\delta}\to y^{\dagger}$ as $\delta\to 0$ is that how to guarantee that $u^{\delta}_{k_{*}(y^{\delta})}$ converges to a solution of (1.2) as $\delta\to 0\mathord{?}$
In general, the answer is no, as stated by Bakushinskii’s veto [3], which asserts that heuristic rules cannot guarantee the convergence in the worst-case scenario for any regularization method. Later, in [9], it is shown that the following assumption guarantee the convergence. We also need this condition in our work.
@@
The family of noisy data $\{y^{\delta}\}$ satisfies the condition $0<\|y^{\delta}-y^{\dagger}\|\rightarrow 0$ as $\delta\rightarrow 0$. Moreover, there exists a constant $\varkappa>0$ such that for every $y^{\delta}$ and for each $u\in S(y^{\delta})$, where $S(y^{\delta}):=\{u^{\delta}_{k}:0\leq k\leq k_{\infty}\}$ is constructed using (3.1), there holds
@@
Under Assumption 3.2, it is immediate to see that Rule 1 comes out with a finite integer $k_{*}$. Indeed, if $k_{\infty}$ is finite, then there is nothing to prove. So we only consider that $k_{\infty}=\infty$. It then follows from Assumption 3.2 that
@@
as $k\rightarrow\infty$. Consequently, we can conclude that a finite integer $k_{*}$ exists, which achieves the minimum of $\Psi(k,y^{\delta})$.
3.2. Analysis for Rule 1
@@
((3.4)) $\Psi(k_{*}(y^{\delta}),y^{\delta})\rightarrow 0\quad\text{as}\quad\delta\rightarrow 0.$
Additionally, we demonstrate that IRSGD iterates $u_{k}^{\delta}$ remains in a ball around $u_{0}$ under certain conditions.
In order to show (3.4), we propose an auxiliary integer $\hat{k}_{\delta}$ defined by the stopping rule
((3.5)) $\|F(u^{\delta}_{\hat{k}_{\delta}})-y^{\delta}\|+\frac{M}{\hat{k}_{\delta}+a}\leq\tau\|y^{\delta}-y^{\dagger}\|<\|F(u^{\delta}_{k})-y^{\delta}\|+\frac{M}{k+a}\quad$
for $0\leq k<\hat{k}_{\delta}$, where $M=\frac{\varrho c(\rho)^{2}\tau}{\Omega(1+\varrho)(1+\eta)},\varrho>0$ is a constant, $\Omega>0,\eta$ is same as in (3.3), $c(\rho)$ is defined in (3.6) and $\tau>1$ is a real number (not to be confused with the one in (1.7)). Adding an extra term in the classical discrepancy principle produces a minor variation of the discrepancy principle (1.7). An essential component in establishing (3.4) is provided by this additional term, which ensures that $\hat{k}_{\delta}\rightarrow\infty$ as $\delta\rightarrow 0$.
@@
Let Assumption 3.1 hold and let $\Omega,\omega$ be positive constants such that $\Omega\geq\omega$. Moreover, let $\kappa\in(0,1)$ and $\varrho>0$ be a sufficiently small constant and
@@
there holds
  • (a)
    the stopping index $\hat{k}_{\delta}$ defined in (3.5) is finite for $\delta>0$ and $\hat{k}_{\delta}\rightarrow\infty$ as $\delta\rightarrow 0.$
  • (b)
    for a solution $\hat{u}$ to (1.2) and for any $0\leq k<\hat{k}_{\delta}$, if $u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$, then $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u})\subset B_{\rho+c(\rho)}(u_{0}).$
@@
First we prove assertion (b). For $k=0$, result trivially follows from (iii) of Assumption 3.1. By assuming that $u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$ for any $0\leq k<\hat{k}_{\delta}$, we demonstrate that $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u}).$ From (3.1), we have
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $=\|u_{k}^{\delta}-\omega_{k}F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0})-\hat{u}\|^{2}$
$=(1-\lambda_{k})^{2}\|u^{\delta}_{k}-\hat{u}\|^{2}+\lambda_{k}^{2}\|u_{0}-\hat{u}\|^{2}+2\lambda_{k}(1-\lambda_{k})\langle u^{\delta}_{k}-\hat{u},u_{0}-\hat{u}\rangle$
((3.9)) $\hskip 14.22636pt+\omega_{k}^{2}\|F^{\prime}_{i_{k}}(u^{\delta}_{k})^{*}(y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}))\|^{2}+2\omega_{k}(1-\lambda_{k})\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),F^{\prime}_{i_{k}}(u^{\delta}_{k})(u^{\delta}_{k}-\hat{u})\rangle$
@@
To estimate 3.2, we can write the last two terms of the sum as
$\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),F_{i_{k}}^{\prime}(u^{\delta}_{k})(u^{\delta}_{k}-\hat{u})\rangle$ $=\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),y_{i_{k}}^{\delta}-y_{i_{k}}^{\dagger}\rangle-\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
((3.10)) $\hskip 5.69054pt-\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),F_{i_{k}}(u^{\delta}_{k})-F_{i_{k}}(\hat{u})-F_{i_{k}}^{\prime}(u^{\delta}_{k})(u^{\delta}_{k}-\hat{u})\rangle,$
@@
where we have used (3.2) and the inequality $2ab\leq a^{2}+b^{2}$ for $a,b\in\mathbb{R}$ in obtaining 3.2.
By incorporating 3.2, 3.2, (3.3) and (3.2) in 3.2, we attain that
@@
$+2\lambda_{k}(1-\lambda_{k})\|u^{\delta}_{k}-\hat{u}\|\|u_{0}-\hat{u}\|$
((3.12)) $+\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|\big{(}2\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|$
$-\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\big{)}.$
@@
Using the fact that $\omega_{k}\in[\omega,\Omega],u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$ and $\hat{u}\in B_{\rho}(u_{0})$ along with inequality (3.13) in 3.2, we have
@@
((3.14)) $-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-2\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
$+\frac{2M\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|}{\tau(k+a)}.$
@@
which is true for $c(\rho)$. From (3.6) along with the inequality $2ab\leq a^{2}+b^{2},\forall\hskip 5.69054pta,b\in\mathbb{R}$ and 3.2, we get
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-2\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
$+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau}\left(\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}+\frac{1}{(k+a)^{2}}\right).$
$\leq-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-(2+M)\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
$+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
As $\omega_{k}\in[\omega,\Omega],0\leq\lambda_{k}\leq\lambda_{\text{max}}<1$ and $1-\lambda_{k}\leq 1$, we further attain
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq-\left((2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})-(2+M)\Omega\tau^{-1}(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
$+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\Omega(1+\eta)}{\tau(k+a)^{2}}.$
Moreover, given the definition of $F$ and the measurability of $u_{k}^{\delta}$ with respect to $\mathcal{F}_{k}$, we can write the last inequality as
$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}|\mathcal{F}_{k}]$ $\leq-\left((2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})-(2+M)\Omega\tau^{-1}(1+\eta)\right)\|y^{\delta}-F(u^{\delta}_{k})\|^{2}$

### v3

@@ 2. Iteratively regularized Stochastic gradient descent in Hilbert spaces
To investigate the convergence behavior of the iterates $u_{k}^{\delta}$ in the IRSGD method, we consider the mean squared norm $\mathbb{E}[\|\cdot\|^{2}]$. In this context, the expectation $\mathbb{E}[\cdot]$ is taken with respect to the filtration $\mathcal{F}_{k}$, which is determined by the sequence of random indices $\{i_{1},\dots,i_{k-1}\}$ (see [20]). We denote by $B_{\rho}(u)$ the closed ball in $U$ of radius $\rho$ centered at $u$, where $u\in U$ and $\rho>0$. Furthermore, let $\mathcal{D}_{\rho}(u^{\dagger})$ represent the set of all solutions to (1.2) within the closed ball $B_{\rho}(u^{\dagger})$, which is formally given by
$\mathcal{D}_{\rho}(u^{\dagger}):=\{u\in B_{\rho}(u^{\dagger}):F(u)=y^{\dagger}\}.$
The remainder of this section is structured as follows: Subsection 2.1 introduces the foundational assumptions required for the convergence analysis and provides a detailed exposition of the IRSGD method, including a heuristic stopping criterion as presented in Algorithm 1. Subsection 2.2 discusses a heuristic stopping rule based on a modified discrepancy principle. In this context, we establish the almost sure finiteness of the stopping index defined by this principle and present several key theoretical results related to the proposed scheme.
@@ 2.1. Assumptions
Our analysis demands the following assumptions.
@@
The following is true for $i\in\{0,1,\ldots,P-1\}$, and $\rho>0$:
@@
    $F:U\rightarrow Y^{P}$ is a continuous operator.
@@
    There exists a constant $0<L\leq\frac{1}{2\omega}$ such that
    ((2.1)) $\|F_{i}^{\prime}(u)\|\leq L\ \text{for all}\ u\in B_{\rho+c(\rho)}(u_{0}).$
    (see Lemma 2.1 for the meaning of $c(\rho)$ and $\omega$.) Moreover, constant $\eta$ exists such that
    ((2.2)) $\|F_{i}(\overline{u})-F_{i}(u)-F_{i}^{\prime}(u)(\overline{u}-u)\|\leq\eta\|F_{i}(\overline{u})-F_{i}(u)\|,\ \ 0<\eta<1$
@@
The conditions in Assumption 2.1, including the tangential cone condition ((2.2)) that quantifies the nonlinearity of the forward operator $F_{i}$, are standard in convergence analyses of iterative regularization for nonlinear inverse problems [8, 16, 22] and have been verified for various settings such as nonlinear integral equations and parameter identification in PDEs [8].
Motivated by [9], we consider the following heuristic stopping rule for the IRSGD method.
Rule 1 (Heuristic).
Let $a\geq 1$ be a fixed constant such that
$\Psi(k,y^{\delta}):=(k+a)\sum_{i=0}^{P-1}\|F_{i}(u^{\delta}_{k})-y_{i}^{\delta}\|^{2}.$
We define $k_{*}:=k_{*}(y^{\delta})$ as the integer satisfying
@@
where $k_{\infty}:=k_{\infty}(y^{\delta})$ denotes the largest integer for which $u^{\delta}_{k}\in\mathfrak{D}(F)$ holds for all $0\leq k<k_{\infty}$.
Using the integer $k_{*}:=k_{*}(y^{\delta})$ defined by Rule 1, we call $u^{\delta}_{k_{*}(y^{\delta})}$ as an approximate solution. This gives rise to the main algorithm of the paper, presented as follows.
Algorithm 1 IRSGD with heuristic stopping
1:Given: $F_{i},y_{i}^{\delta}$ for all $i\in\{0,1,\ldots,P-1\}$ and a constant $a\geq 1$
2:Initialize: $u_{0}^{\delta}=u_{0}$, set $k=0$
3:repeat
4:  Choose $\omega_{k}\in[\omega,\Omega]$ and the weight parameter $\lambda_{k}$ such that
$0\leq\lambda_{k}\leq\lambda_{\text{max}}<\frac{1}{2}.$
5:  Randomly pick an index $i_{k}$ from $\{0,1,\ldots,P-1\}$ uniformly.
6:  Update
$u_{k+1}^{\delta}=u_{k}^{\delta}-\omega_{k}F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0}),$
where $F^{\prime}_{i_{k}}$ is the Fréchet derivative of $F_{i_{k}}$, and $u_{0}$ is the initial guess.
7:  Set $k\leftarrow k+1$
8:until
$\Psi(k,y^{\delta}):=(k+a)\sum_{i=0}^{P-1}\|F_{i}(u^{\delta}_{k})-y_{i}^{\delta}\|^{2}$
does not achieve its minimum (for the details see heuristic rule 1)
9:return  $u^{\delta}_{k_{*}(y^{\delta})}$
@@
Rule 1 is straightforward to implement in practice, as detailed in Section 4. Throughout the iterative procedure, the value of $\Psi(k,y^{\delta})$ is monitored as a function of $k$. After a sufficient number of iterations, the process is terminated and $k_{*}$ is selected that minimizes $\Psi(k,y^{\delta})$. However, an important challenge associated with the nonlinear Landweber iteration [12] lies in determining an appropriate upper bound on the number of iterations. This is particularly critical due to its local convergence properties, which can lead to premature stopping or inaccurate identification of $k_{*}$. Therefore, the limit on the upper bound must be sufficiently large to avoid inaccurately estimating $k_{*}$ at the first local minimum, which is often not the global minimum [12].
Remark 2.2.
Evaluating the residual $\|F(u_{k}^{\delta})-y^{\delta}\|^{2}=\sum_{i=0}^{P-1}\|F_{i}(u_{k}^{\delta})-y_{i}^{\delta}\|^{2}$ at every iteration of Algorithm 1 for heuristic rule 1 can be computationally expensive. A practical alternative is to perform this computation at prescribed intervals, following the approach used in stochastic gradient methods [13, 21], or to exploit randomized SVD techniques in problems with intrinsic low-rank structure [23]. Similar computational overheads have been reported for the discrepancy principle [6, 11]. Mitigating these costs in the heuristic setting, potentially by adapting the strategy in [11], is a promising direction for future research.
Algorithm 1 leads to the question of how to ensure that $u^{\delta}_{k_{*}(y^{\delta})}$ converges to a solution of (1.2) as $\delta\to 0.$ Specifically, we seek to determine the conditions under which this convergence holds for a sequence of noisy data $\{y^{\delta}\}$, where $y^{\delta}\to y^{\dagger}$ as $\delta\to 0$.
In general, the answer is inadequate, as stated by Bakushinskii’s veto [2], which asserts that heuristic rules cannot guarantee convergence in the worst-case scenario for any regularization method. Subsequently, in [9], a noise condition was introduced for linear problems in Hilbert spaces. This condition was later generalized under the Hanke-Raus rule (see Assumption 2.2) to examine the convergence of nonlinear regularization methods as well as iterative regularization methods in Banach spaces [18, 19, 32, 36]. We also utilize this condition in our work.
@@
The family of noisy data $\{y^{\delta}\}$ satisfies the condition $\|y^{\delta}-y^{\dagger}\|\rightarrow 0$ as $\delta\rightarrow 0$. Moreover, there exists a constant $\varkappa>0$ such that for every $y^{\delta}$ and for each $u\in S(y^{\delta})$, where $S(y^{\delta}):=\{u^{\delta}_{k}:0\leq k\leq k_{\infty}\}$ is constructed using (1.11), there holds
@@
It directly follows from Assumption 2.2 that Rule 1 yields a finite integer $k_{*}$. Indeed, if $k_{\infty}$ is finite, then $k_{*}$ is trivially finite. Therefore, we assume $k_{\infty}=\infty$. From Assumption 2.2 there holds
@@
as $k\rightarrow\infty$. Consequently, a finite integer $k_{*}$ exists that achieves the minimum of $\Psi(k,y^{\delta})$.
Remark 2.3.
It is worth emphasizing that Assumption 2.2, while abstract and difficult to verify analytically, currently offers the most practical framework for theoretical analysis, as also noted in [32]. Its validity can, at present, only be supported through numerical experiments. A detailed numerical investigation of heuristic parameter choice rules, beyond the Hanke–Raus rule, for the nonlinear Landweber iteration is provided in [12].
2.2. Analysis for Heuristic Rule 1
@@
((2.3)) $\mathbb{E}\left[\Psi(k_{*}(y^{\delta}),y^{\delta})\right]\rightarrow 0\quad\text{as}\quad\delta\rightarrow 0.$
We show the same in Lemma 2.2. Additionally, we demonstrate that IRSGD iterates $u_{k}^{\delta}$ remains in a ball around $u_{0}$ under certain conditions.
To establish ((2.3)), we introduce an auxiliary index $\hat{k}_{\delta}$, determined by the following stopping criterion.
Rule 2 (Modified discrepancy).
The stopping index $\hat{k}_{\delta}$ is defined as the smallest integer satisfying
((2.4)) $\|F_{i}(u^{\delta}_{\hat{k}_{\delta}})-y_{i}^{\delta}\|+\frac{M}{\hat{k}_{\delta}+a}\leq\tau\|y_{i}^{\delta}-y_{i}^{\dagger}\|<\|F_{i}(u^{\delta}_{k})-y_{i}^{\delta}\|+\frac{M}{k+a},\quad\forall i=0,1,\ldots,P-1,$
for all $0\leq k<\hat{k}_{\delta}$, where
$M=\frac{\varrho\,c(\rho)^{2}\,\tau}{\Omega\,(1+\varrho)\,(1+\eta)},$
with $\varrho>0$, $\Omega>0$, $\eta$, $\tau>1$, and $c(\rho)$ defined in ((2.2)), ((1.7)), and ((2.5)), respectively.
This modified discrepancy principle represents a slight augmentation of the classical version given in ((1.7)), achieved by incorporating an additional penalization term. This term plays a pivotal role in the convergence analysis, particularly in establishing the result in ((2.3)), as it guarantees that $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$.
Remark 2.4.
It is important to emphasize that both the stopping indices $k_{*}$ and $\hat{k}_{\delta}$ depend on the iterates $u_{k}^{\delta}$, which are themselves random variables. Consequently, $k_{*}$ and $\hat{k}_{\delta}$ are also random quantities, adding a substantial layer of complexity to the theoretical analysis compared to the deterministic settings considered in [18, 19, 32].
[…]
Algorithm 2 IRSGD with modified discrepancy stopping
1:Given: operators $F_{i}$ and noisy data $y_{i}^{\delta}$ for $i=0,1,\dots,P-1$; constant $a\geq 1$; parameters $\omega,\Omega>0$, $\lambda_{\max}\in[0,\tfrac{1}{2})$; discrepancy parameters $\tau>1$ and $M>0$ (as in ((2.4))).
2:Initialize: $u_{0}^{\delta}=u_{0}$ (initial guess), set $k\leftarrow 0$.
3:repeat
4:  Choose $\omega_{k}\in[\omega,\Omega]$ and $\lambda_{k}$ with $0\leq\lambda_{k}\leq\lambda_{\max}$.
5:  Select index $i_{k}$ uniformly at random from $\{0,1,\dots,P-1\}$.
6:  Compute update
$u_{k+1}^{\delta}\;=\;u_{k}^{\delta}-\omega_{k}\,F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}\big{(}F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\big{)}-\lambda_{k}\,(u_{k}^{\delta}-u_{0}).$
7:  Set $k\leftarrow k+1$.
8:until the modified-discrepancy condition ((2.4)) is satisfied for the current iterate, i.e., the index $\hat{k}_{\delta}$ defined by
$\|F_{i}(u^{\delta}_{\hat{k}_{\delta}})-y_{i}^{\delta}\|+\frac{M}{\hat{k}_{\delta}+a}\leq\tau\|y_{i}^{\delta}-y_{i}^{\dagger}\|\quad\text{for all }i=0,1,\dots,P-1,$
holds (see Rule 2).
9:return  $u^{\delta}_{\hat{k}_{\delta}}$
@@
Let Assumption 2.1 hold and let $\Omega,\omega$ be positive constants such that $\Omega\geq\omega_{k}\geq\omega$. Moreover, let $\kappa\in(0,1)$ and $\varrho>0$ be a sufficiently small constant (see Remark 2.5) and
@@
[…]
@@
[…]
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $=(1-\lambda_{k})^{2}\|u^{\delta}_{k}-\hat{u}\|^{2}+\lambda_{k}^{2}\|u_{0}-\hat{u}\|^{2}+2\lambda_{k}(1-\lambda_{k})\langle u^{\delta}_{k}-\hat{u},u_{0}-\hat{u}\rangle$
[…]
@@
To estimate 2.2, we can write the last two terms of the sum as
[…]
@@
where we have used (2.1) and the inequality $2ab\leq a^{2}+b^{2}$ for $a,b\in\mathbb{R}$ in obtaining 2.2.
By incorporating (2.1), (2.2), 2.2, and 2.2 in 2.2, we attain that
@@
$\quad+2\lambda_{k}(1-\lambda_{k})\|u^{\delta}_{k}-\hat{u}\|\|u_{0}-\hat{u}\|$
((2.11)) $\quad+\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|\big{(}2\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|$
$\quad-\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\big{)}.$
@@
[…]
@@
((2.13)) $\quad-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-2\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
$\quad+\frac{2M\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|}{\tau(k+a)}.$
@@
which is true for $c(\rho)$. From (2.5) along with the inequality $ab\leq a^{2}+b^{2},\forall\hskip 5.69054pta,b\in\mathbb{R}, 2.2$, $\omega_{k}\in[\omega,\Omega],0\leq\lambda_{k}\leq\lambda_{\text{max}}<1$ and $1-\lambda_{k}\leq 1$, we get
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-(2+M)\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
$\quad+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}$
$\leq-\left((2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})-(2+M)\Omega\tau^{-1}(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$